OpenStudy (anonymous):

find the derivative of f(x)= sqrt(4-x^2)

OpenStudy (anonymous):

(4-x^2)^1/2 1/2(4-x^2)^-1/2 1/2sqrt(4-x^2) * (-2)

OpenStudy (anonymous):

sorry , it's * (-2x) ^^

OpenStudy (haseeb96):

f(x)=sqrt(4-x^2) differentiate w.r.t. x df(x)\dx=d{sqrt4-x^2}\dx =(1)(4-x^2)\2 * 0 - 2x = (4-x^2)(-2x)\2 = -4x + 2x^3\2 df(x)\dx = 2x^3-4x\2 this is the answer

OpenStudy (anonymous):

@Haseeb96 why did you multiple the df(x)/dx by df(x)?

OpenStudy (anonymous):

I have a question if you have two inequalities and one is a dotted line on the y value of -2 and the other is a solid line is on the .5 x value and the shaded region is quadrant one and part of quadrant four would the point of intersection not be a solution or be a solution. if someone could help that would be great.

OpenStudy (ranga):

f(x)= sqrt(4-x^2) f(x) = (4 - x^2)^(1/2) Use the power rule and then the chain rule. f'(x) = 1/2 * (4 - x^2)^(-1/2) * (-2x) = -x * (4 - x^2)^(-1/2) = -x / (4 - x^2)^(1/2) = \[\Large \frac{ -x }{ \sqrt{4 - x^2}}\]